The height of a golf ball after being hit is approximated by the equation h=18t-4t^2, where h is the height above the ground in meters t seconds after the ball is hit. Algebraically determine the approximate instantaneous rate of change of height of the golf ball when the time is 3.0 seconds, and describe what is happening to the ball at that instant. Answer by stanbon(75887) 
 The height of a golf ball after being hit is approximated by the equation h=18t-4t^2, where h is the height above the ground in meters t seconds after the ball is hit. Algebraically determine the approximate instantaneous rate of change of height of the golf ball when the time is 3.0 seconds, and describe what is happening to the ball at that instant.
------------------
Take the derivative of h(t).
h'(t) = 18-8t
-------
Find h'(3) = 18-24 = -6 (This is the rate of change at t=3 seconds)
================
Cheers,
Stan H.